OpenStudy (loser66):

Let \(f_1, f_2,f_3\) be continuously differentiable functions from \(\mathbb R^4\) to \(\mathbb R\). Give suffice conditions so that the equations \(f_1(x,y,z, t)=0\\f_2(x,y,z,t)=0\\f_3(x,y,z,t)=0\) can be solved for x,y,z in term of t. Please, help

OpenStudy (loser66):

@SithsAndGiggles @amistre64 @phi @perl

OpenStudy (amistre64):

i wouldnt even know where to start with this at the moment. srry

OpenStudy (loser66):

It's Ok. Thanks anyway. :)

OpenStudy (nincompoop):

http://www.owlnet.rice.edu/~fjones/chap2.pdf

OpenStudy (loser66):

@nincompoop Thanks but it is not the same as my problem. :)

OpenStudy (loser66):

My attempt: Let \(f_1,f_2,f_3:\mathbb R^4 \rightarrow \mathbb R\) . Let \(f:\mathbb R^4\rightarrow \mathbb R^3\), \(f=(f_1,f_2,f_3)\) . Hence f is smooth function

OpenStudy (loser66):

The derivative matrix of f is \[D(f)=\left[\begin{matrix}\dfrac{\partial f_1}{\partial x}&\dfrac{\partial f_1}{\partial y}&\dfrac{\partial f_1}{\partial z}\\\dfrac{\partial f_2}{\partial x}&\dfrac{\partial f_2}{\partial y}&\dfrac{\partial f_2}{\partial z}\\\dfrac{\partial f_3}{\partial x}&\dfrac{\partial f_3}{\partial y}&\dfrac{\partial f_3}{\partial z}\end{matrix}\right]\]

OpenStudy (loser66):

But I don't know where to go from this. :(

OpenStudy (loser66):

How about this @rational @zzr0ck3r

OpenStudy (zzr0ck3r):

I have no idea either. I really hate analysis and much prefer topology.

OpenStudy (loser66):

I will take topology next semester, is it hard?? @zzr0ck3r

OpenStudy (zzr0ck3r):

I don't think so. It is much more general.

OpenStudy (zzr0ck3r):

I find it makes analysis easier, but we are not this far yet.

OpenStudy (loser66):

any recommendation for preparing for the course?

OpenStudy (zzr0ck3r):

Munkres Topology

OpenStudy (zzr0ck3r):

it's a staple

OpenStudy (loser66):

are they books? any site?

OpenStudy (zzr0ck3r):

yeah it's a book. I don't really watch videos of lectures, I learn much more from the books.

OpenStudy (zzr0ck3r):

You can readily find it if you google

OpenStudy (zzr0ck3r):

for free even! The book is super old.

OpenStudy (loser66):

I will, thanks for the tips. :)

OpenStudy (zzr0ck3r):

If you have any questions next term, tag me:) I love that stuff